OpenStudy (anonymous):

Two of the sides of a triangular lot have lengths of 90 feet and 80 feet. If the angle opposite the 80-foot side is 59°, what is the length of the third side of the lot to the nearest tenth of a foot? A. 10 ft B. 65.1 ft C. 67.5 ft D. 57.9 ft

OpenStudy (anonymous):

67.5!

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

You need to apply the Sine Rule twice, firstly to get te angle opposite the 90 foot length. Call this angle A. Now Sin A/90 = sin 59/80 which gives Sin A = 90 x Sin 59/80 hence A = 74.647.... So the third angle in the triangle is 180 - 59 - 74.647... = 46.3529.... Call the third length c. Now c/Sin 46.3529... = 80/ Sin 59 so c = 67.53455... 67.5 is the answer.